OpenStudy (anonymous):

Alvin can go upstream at a rate of 4.8 miles per hour and downstream at a rate of 8.5 miles per hour. One day while Alvin was on the river, he received a call telling him he need to turn around and come home. So he turned around and went back to his starting point. If his entire trip took 17 hours, how far did he travel on the river?

OpenStudy (mathstudent55):

Distance each way: d upstream: speed 4.8 mph, time t1 downstream: speed 8.5 mph, time t2 Total time: 17 h distance = speed * time upstream d = 4.8 * t1 downstream d = 8.5 * t2 t1 + t2 = 17 t1 = 17 - t2 d = 4.8(17 - t2) d = 8.5(t2) 4.8(17 - t2) = 8.5t2 81.6 - 4.8t2 = 8.5t2 81.6 = 13.3t2 t2 = 6.135 h d = 8.5 * t2 = 8.5 * 6.135 = 52.15 miles

OpenStudy (anonymous):

thank you this really helped i get what you did and i was able to do two other similar questions

OpenStudy (mathstudent55):

Great. you're welcome.

OpenStudy (anonymous):

Refer to a solution using the Mathematica computer program.